Accelerator Update for February 2, 2009 - February 4, 2009 |
Monday Feb 2 The day shift began with Operations monitoring the Tevatron (TeV) store 6766 (120.46E30) and the Recycler stash (154.0E10), with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beams, and with the Linac in access to repair B:CHOP. B:CHOP repairs were completed at 9:00 AM. Beam returned to Pbar, NuMI and MiniBooNE at 9:17 AM. Operations terminated store 6766 at 5:11 PM. CDF and D0 began a controlled access at 5:28 PM. D0 came out of access at 6:03 PM. CDF came out of access at 7:04 PM. Operations put the TeV into shot setup at 9:27 PM. Operations established store 6770 at 10:47 PM with an initial luminosity of 316.44E30. Tuesday Feb 3 The midnight shift began with Operations monitoring store 6770 (227.54E30) and the stash (85.5E10), with Pbar stacking (16.5E10), and with NuMI and MiniBooNE taking beam. Operations terminated store 6770 at 2:09 PM. Operations put the TeV into shot setup at 2:21 PM. Operations established store 6771 at 4:29 PM with an initial luminosity of 321.31E30. At 7:19 PM, Operations started to experience problems with the Booster Chopper (B:CHOP), experts were called to investigate. No beam for stacking or to the users. B:CHOP repairs were completed at 10:12 PM. Beam returned to Pbar, NuMI and MiniBooNE at 10:21 PM. Wednesday Feb 4 The midnight shift began with Operations monitoring store 6771 (117.33E30) and the stash (143.6E10), with Pbar stacking (27.6E10), and with NuMI and MiniBooNE taking beam. The Plans for today The plans are to stack, stash, and store. |
|
More Information Comments and Suggestions |
| last modified 02/04/2009 email Fermilab |
| Security, Privacy, Legal | |